OpenStudy (anonymous):

Determine the value of k so that one root of x^2 + 8x + k = 0 is twice the other root.

myininaya (myininaya):

Say the roots are \[x=r_1, r_2\] We are given that one of the roots is twice the other. So we have this: \[x=r_1\] \[x=r_2=2r_1\] Ok so we have \[x=r_1 => x-r_1 \text{ is a factor}\] \[x=2r_1 => x-2r_1 \text{ is another factor}\] so when we factor \[x^2+8x+k\] we are to get \[(x-2r_1)(x-r_1)\] So what happens if we expand this (multiply these factors) We get \[(x-2r_1)(x-r_1)=x^2-r_1x-2r_1x+2r_1^2=x^2+x(-r_1-2r_1)+2r_1^2\] But this is suppose to be equal to \[x^2+8x+k\] So thats mean you have \[-r_1-2r_1=8 \text{ and } 2r_1^2=k\]

myininaya (myininaya):

\[-3r_1=8 => r_1=\frac{8}{-3}\] \[2r_1^2=k => r_1=\pm \sqrt{\frac{k}{2}}\] => \[\frac{8}{-3}=\pm \sqrt{\frac{k}{2}}\]

myininaya (myininaya):

so we can say solve this for k

myininaya (myininaya):

square both sides

myininaya (myininaya):

\[\frac{64}{9}=\frac{k}{2}\]

myininaya (myininaya):

not multiply both sides by 2

myininaya (myininaya):

now*

myininaya (myininaya):

\[\frac{128}{9}=k\]

OpenStudy (anonymous):

i kept on messing up

OpenStudy (anonymous):

(x-a) (x-2a) = x^2 - 3ax + 2a^2 x^2 + 8x + k 8 = -3a a = 8/-3 k =128/9

myininaya (myininaya):

yay! :)

OpenStudy (anonymous):

@myininaya: Why don't you use vieta's formula? Setting \( \alpha \) and \( 2 \alpha \) are the roots of the equation, then $$ \alpha + 2 \alpha = -8 \Rightarrow \alpha = -\frac 83 $$ and, $$ 2 \alpha^2 = k \Rightarrow k = \frac{ 128} {9} $$ and we are done!

myininaya (myininaya):

i like my way because its longer lol

myininaya (myininaya):

very nice fool

OpenStudy (anonymous):

Thanks :D Simple is beautiful :)